### v2

@@
[…]
tcb@breakable marginparsep has been altered.
topmargin has been altered.
marginparpush has been altered.
The page layout violates the ICML style. Please do not change the page layout, or include packages like geometry, savetrees, or fullpage, which change it for you. We’re not able to reliably undo arbitrary changes to the style. Please remove the offending package(s), or layout-changing commands and try again.
Training of Scaffolded Language Models
with Language Supervision: A Survey
Matthieu Lin 1,2 Jenny Sheng 2 Andrew Zhao 1
Shenzhi Wang 1 Yang Yue 1 Victor Shea-Jay Huang 3 Huan Liu 4 Jun Liu 4
[…]
${}^{1\,}$eap Lab, Tsinghua University ${}^{2\,}$, Tsinghua University ${}^{3\,}$hanghai AI Lab ${}^{4\,}$i’an Jiaotong University
linmatthieu@gmail.com, jsheng415@outlook.com, gaohuang@tsinghua.edu.cn, liuyongjin@tsinghua.edu.cn
footnotemark: footnotetext: ✉ Corresponding author(s)
This survey organizes the intricate literature on the design and optimization of emerging structures around post-trained LMs. We refer to this overarching structure as scaffolded LMs and focus on LMs that are integrated into multi-step processes with tools. We view scaffolded LMs as semi-parametric models wherein we train non-parametric variables, including the prompt, tools, and scaffold’s code. In particular, they interpret instructions, use tools, and receive feedback all in language. Recent works use an LM as an optimizer to interpret language supervision and update non-parametric variables according to intricate objectives. In this survey, we refer to this paradigm as training of scaffolded LMs with language supervision. A key feature of non-parametric training is the ability to learn from language. Parametric training excels in learning from demonstration (supervised learning), exploration (reinforcement learning), or observations (unsupervised learning), using well-defined loss functions. Language-based optimization enables rich, interpretable, and expressive objectives, while mitigating issues like catastrophic forgetting and supporting compatibility with closed-source models. Furthermore, agents are increasingly deployed as co-workers in real-world applications—such as Copilot in Office tools or software development. In these mixed-autonomy settings, where control and decision-making are shared between human and AI, users point out errors or suggest corrections. Accordingly, we discuss agents that continuously improve by learning from this real-time, language-based feedback and refer to this setting as streaming learning from language supervision.
@@ 1 Introduction
Figure 1: Non-parametric training of scaffolded LMs focuses on learning from language. Parametric training has excelled in learning from demonstration (supervised learning), exploration (reinforcement learning), and observation (unsupervised learning) by using well-defined loss functions. Instead, non-parametric training enables using a LM to optimize in the language space, allowing for efficient and interpretable learning with fuzzy objectives and rich textual feedback. As scaffolded LMs interact with digital systems operated by humans, mixed autonomy Cui et al. (2023) becomes increasingly important. We anticipate that learning from language will serve as a crucial framework for scaffolded LMs that learn across episodes Silver & Sutton (2025) (streaming learning) from rich user feedback in mixed autonomy. Importantly, it circumvents catastrophic forgetting Hadsell et al. (2020), which prevents parametric training from learning across episodes. We refer readers to Sec. 4.4 for further discussion.
@@
Scaffolded language models embed language models (LMs) in a framework that extends their ability beyond traditional NLP tasks to perform open-ended digital automation tasks. Notably, they mark the beginning of a new type of programs for digital automation OpenAI (2018). It uses language to interface with humans, digital systems, and developer-defined code. This opens up the opportunity for mixed autonomy where humans and AI share control and decision-making responsibilities to achieve a common goal Wang et al. (2024c); Cui et al. (2023).
@@
We define a scaffolded LM as a structure that integrates the LM into a multi-step process with tools. Specifically, the post-training interface of LM enables the developer to specify the system behavior in natural language Grattafiori et al. (2024). The developer specifies available tools, and the LM executes them by generating the function calls and input arguments. Tools Wang et al. (2024e) typically correspond to digital systems or developer-defined code. The scaffold Wu et al. (2023); Liu (2022); AI (2024); Khattab et al. (2023) refers to a surrounding structured code that manages control flow (multi-step process), maintains state (prompt to LM), and executes tools based on responses of the LM. Thus, the resulting scaffolded LM can be viewed as a semi-parametric model, where the LM is the parametric component. In particular, the scaffold enables an interpretable and efficient way to modify the behavior of the ‘program’. For instance, we can add new or modify tools corresponding to APIs to interact with the digital world, developer-defined code to perform computations, or even other entities such as the human or another (scaffolded) LM. As well as creating detailed prompts with intricate constraints for the LM. The design and optimization of scaffolded LMs requires significant trial and error by expert developers. This involves identifying the LM’s failure modes by running the scaffolded LM on examples of user queries and then addressing them one by one by refining the system prompt, input template, available tools, or the scaffold’s code itself.
@@
Accordingly, recent works Yuksekgonul et al. (2025); Zhao et al. (2024a) have proposed to use an LM as an optimizer that interprets intricate objectives, execution traces, and rich execution feedback to generate updated non-parametric variables. In practice, scaffolded LMs are naturally exposed to a wide variety of rich language feedback, such as user feedback or execution errors, that reveal about the failure modes. Beyond learning from execution feedback, the LM optimizer also extracts insights or reusable tools from successful execution traces Wang et al. (2025c); Zhao et al. (2024a).
@@
As shown in Fig. 1, training of non-parametric variables focuses on learning from language. Notably, parametric training excels in learning from demonstrations (supervised learning) Ouyang et al. (2022), exploration (reinforcement learning) DeepSeek-AI et al. (2025) , and observations (unsupervised learning)Brown et al. (2020), using well-defined loss functions LeCun et al. (2015). In contrast, optimization in the language space Cheng et al. (2023) provides a learning system with intricate objectives expressed in natural language that is efficient and interpretable. The training of non-parametric variables circumvents key limitations of gradient descent, such as catastrophic forgetting Hadsell et al. (2020), while being compatible with closed-source models Anthropic (2024b); OpenAI (2024). In this survey, we dub this framework training of scaffolded language model with language supervision.
@@
This survey organizes the intricate literature of compound AI systems Zaharia et al. (2024), LM pipelines Khattab et al. (2023), prompt optimization Zhou et al. (2023), language agents OpenAI (2025); Yao (2024); Xi et al. (2023); Masterman et al. (2024), AI workflow optimization Cheng et al. (2024); Yuksekgonul et al. (2025); Yin (2024), and experiential learning agents Zhao et al. (2024a) under the concept of training of scaffolded language model with language supervision. It focuses on the multi-step aspect that involves optimization beyond the prompt of a single LLM call Zhou et al. (2023) to multiple LM calls and tools. Specifically, compound AI systems Zaharia et al. (2024); Lin et al. (2024) focus on the systemic perspective (the interconnection of components) and overlook the multi-step aspect. Meanwhile, the literature in language agents Xi et al. (2023) neglects the efficiency of LM pipelines (also called workflows) Khattab et al. (2023) that leverage the task structure to reduce the burden in decision-making and supervise the LM with clearer instructions.
To help readers navigate through the paper, we briefly introduce the organization of this survey. As shown in Fig. 2, The survey contains three sections.
@@
  • Sec. 2 focuses on how the post-training interface of LMs interacts with the scaffold. We identify two types of scaffold: agent and workflow. And focus on important design aspect of agent and workflow. Finally, we reflect on the trend of recent benchmarks.
  • Then, Sec. 3 organizes the intricate literature that uses a LM as an optimizer to train the non parametric variables. We identify a trend whereby these methods focus on training more complex scaffold (where complexity is defined by how much heavy lifting is done by the scaffold v.s. the LM to solve the task).
  • Finally, our discussion in Sec. 4 points out a key capability missing from current scaffolded LM: the ability to continuously learn across episodes Silver & Sutton (2025); Yao (2025). This motivates a novel setting for training scaffolded LMs with language supervision. This discussion introduces recent progress in RL for LMs as well as post-training. It grounds our understanding of non-parametric training into the broader field. It highlights how current applications of non-parametric training focuses on an inappropriate setting, where RL from verifiable reward Lambert et al. (2025) is a much better alternative.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=10pt, l sep=10pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center , where level=1text width=14em,font=,, where level=2text width=14em,font=,, where level=3text width=10em,font=,, [ Training of Scaffolded Language Models With Language Supervision, ver, color=mypink!75!black, fill=mypink!100, text=black [ Scaffolded Language Models §2, color=myblue!75!black, fill=myblue!100, text=black [ Post-Trained
@@
Language Models §2.1, color=myblue!75!black, fill=myblue!60, text=black ] [ Tools §2.2, color=myblue!75!black, fill=myblue!60, text=black ] [ Multi-step Process §2.3, color=myblue!75!black, fill=myblue!60, text=black [ Agent §2.3.1, color=myblue!75!black, fill=myblue!60, text=black ] [ Workflow §2.3.2, color=myblue!75!black, fill=myblue!60, text=black ] ] [ Benchmarks §2.4, color=myblue!75!black, fill=myblue!60, text=black ] ] [ Training with Language Supervision §3, color=myyellow!75!black, fill=myyellow!100, text=black [ Prompt Optimization §3.1, color=myyellow!75!black, fill=myyellow!60, text=black ] [ Experiential Learning §3.2, color=myyellow!75!black, fill=myyellow!60, text=black [ Insights §3.2.1, color=myyellow!75!black, fill=myyellow!60, text=black ] [ Tools §3.2.2,color=myyellow!75!black, fill=myyellow!60, text=black ] ] [ AutoDiff frameworks §3.3, color=myyellow!75!black, fill=myyellow!60, text=black [ Graph Abstraction §3.3.1, color=myyellow!75!black, fill=myyellow!60, text=black ] [ Graph Execution and
Optimization §3.3.2,color=myyellow!75!black, fill=myyellow!60, text=black ] ] ] [ Discussion §4, color=mygreen!75!black, fill=mygreen!100, text=black [ Post-Training §4.1, color=mygreen!75!black, fill=mygreen!60, text=black ] [ Learned LM Optimizer §4.2, color=mygreen!75!black, fill=mygreen!60, text=black ] [ Multi-Turn RL §4.3, color=mygreen!75!black, fill=mygreen!60, text=black ] [ Streaming Learning
with Language Supervision§4.4, color=mygreen!75!black, fill=mygreen!60, text=black ] ] ]
@@
Figure 2: Organization of this survey. This survey introduces a unifying paradigm centered on learning from language using the semi-parametric capabilities of scaffolded language models. Despite the complexity and fragmentation of the literature across domains such as language agents, prompt optimization, experiential learning, and compound AI systems, we adopt a deliberately simple and clear organizational structure. This simplicity is not a limitation, but a reflection of our effort to distill a coherent framework from a rapidly evolving field. It enables us to outline key research opportunities in learning from language supervision.
@@ 2 Scaffolded Language Model
A scaffolded LM refers to an LM embedded in a multi-step process with tools. Such a structure allows for extending an LM’s utility to interact with digital systems in similar ways as humans do Wang et al. (2024c). While prior works have considered a meta-generation algorithm as an agent and employed an LM optimizer to generate its design Hu et al. (2024); Zhang et al. (2025); Saad-Falcon et al. (2024), our definition of a scaffolded LM does not include meta-generation techniques Welleck et al. (2024) such as best-of-n sampling or tree of thought as a scaffold. Meta-generation refers to search algorithms Silver et al. (2016) for reasoning tasks Cobbe et al. (2021) where each step consists of a call to an LM. Importantly, in scaffolded LMs, each call to a LM expects a tool response.
To provide readers with a practical understanding of how this structure allows LMs to perform sophisticated cognitive tasks and improve from non-parametric training, we organize this section as follows. We introduce the post-training interface of language models that allows interaction with tools and defines their behavior in natural language. Next, we define what we refer to as a tool. Furthermore, we show how the multi-step process can be controlled to implement priors into how the task is solved. Finally, we discuss benchmarks to understand the shift induced by scaffolded LMs. Specifically, we are excited by novel applications where humans share control and decision-making responsibilities with AI Windsurf .

### v3

@@
Abstract
This survey organizes the intricate literature on the design and optimization of emerging structures around post-trained LMs. We refer to this overarching structure as scaffolded LMs and focus on a scaffold that integrates LMs in a multi-step process with tools. We view scaffolded LMs as semi-parametric models wherein we optimize prompts and tools, which we refer to as non-parametric variables. In particular, scaffolded LMs interpret instruction, use tools, and receive feedback all in language. Recent works use an LM as an optimizer to interpret language supervision and update non-parametric variables according to intricate objectives. In this survey, we refer to this paradigm as training of scaffolded LMs with language supervision. A key feature of non-parametric training is the ability to learn from language. Parametric training excels in learning from demonstration (supervised learning), exploration (reinforcement learning), or observations (unsupervised learning), using well-defined loss functions. Optimization in the space of language enables rich, interpretable, and expressive objectives, while mitigating issues like catastrophic forgetting and supporting compatibility with closed-source models. Furthermore, agents are increasingly deployed as co-workers in real-world applications—such as Copilot in Office tools or software development. In these mixed-autonomy settings, where control and decision-making are shared between humans and AI, users provide feedback by identifying errors or offering corrections. Accordingly, we discuss agents that inhabit streams of experience by learning from this language-based feedback.
keywords:
Compound AI Systems, Language Agent, Mixed-Autonomy, Prompt Optimization, Experiential Learning Agents
@@ 1 Introduction
Figure 1: Non-parametric training of scaffolded LMs focuses on learning from language. Parametric training has excelled in learning from demonstration (supervised learning), exploration (reinforcement learning), and observation (unsupervised learning) by using well-defined loss functions. Instead, non-parametric training enables using an LM to optimize in the language space, allowing for efficient and interpretable learning with fuzzy objectives and rich textual feedback. As scaffolded LMs interact with digital systems operated by humans, mixed autonomy becomes increasingly important [1]. We anticipate that learning from language will serve as a crucial framework for scaffolded LMs that inhabit streams of experience from rich user feedback in mixed autonomy [2]. Importantly, it circumvents catastrophic forgetting, supports compatibility with closed-source models, and is interpretable [3]. We refer readers to Sec. 5 for further discussion.
@@
Scaffolded language models embed language models (LMs) in a framework that extends their ability beyond traditional NLP tasks to perform open-ended digital automation tasks. Notably, they mark the beginning of a new type of programs for digital automation [4]. It uses language as the interface between humans and digital systems. This opens up the opportunity for mixed autonomy where humans and AI share control and decision-making responsibilities for digital tasks [5, 1].
@@
Following the terminology from OpenAI [6], we define a scaffolded LM as a structure that integrates the LM into a multi-step process with tools. Specifically, the post-training interface of the LM enables the developer to specify the system behavior in natural language. The developer specifies available tools, and the LM executes them by generating the function calls and input arguments. Tools [7] correspond to external capabilities the model can invoke (see the glossary in Fig. 3). The scaffold refers to a surrounding structured code that manages control flow (multi-step process), maintains state (chat history), and executes tools based on responses of the LM. Thus, the resulting scaffolded LM can be viewed as a semi-parametric model, where the LM is the parametric component and non-parametric components correspond to tools and prompts. In particular, non-parametric components enables an interpretable and efficient way to modify the behavior of the ‘program’. For instance, we can add new or modify tools corresponding to APIs to interact with the digital world, developer-defined code to perform computations, or even other entities such as the human or another (scaffolded) LM [8]. As well as creating detailed prompts with intricate constraints. The optimization of scaffolded LMs requires significant trial and error by expert developers. This involves identifying the LM’s failure modes by running the scaffolded LM on examples of user queries and then addressing them one by one by refining the system prompt, input template, available tools, or the scaffold’s code itself.
@@
Accordingly, recent works [9, 10] have proposed to use an LM as an optimizer that interprets intricate objectives, execution traces, and rich execution feedback to generate updated non-parametric variables. In practice, scaffolded LMs are naturally exposed to a wide variety of rich language feedback, such as user feedback or execution errors, that reveal about the failure modes. Beyond learning from execution feedback, the LM optimizer can also extract insights or reusable tools from successful execution traces [11, 10].
@@
As shown in Fig. 1, training of non-parametric variables focuses on learning from language. Notably, parametric training excels in learning from demonstrations (supervised learning) [12], exploration (reinforcement learning) [13] , and observations (unsupervised learning)[14], using well-defined loss functions [15]. In contrast, optimization in the language space [16] provides a learning system with intricate objectives expressed in natural language that is efficient and interpretable. The training of non-parametric variables circumvents key limitations of gradient descent, such as catastrophic forgetting [3], while being compatible with closed-source models [17, 18]. In this survey, we refer to this framework as training of scaffolded language models with language supervision.
@@
This survey presents a taxonomy for understanding this paradigm shift from parametric training to non-parametric training. In particular, it organizes the intricate literature on compound AI systems [19], LM pipelines [20], prompt optimization [21, 22], language agents [23, 24, 25, 26], AI workflow optimization [27, 9, 28], and experiential learning agents [10] under the concept of training of scaffolded language models with language supervision. It focuses on the multi-step aspect that involves optimization beyond the prompt of a single LLM call [21] to multiple LM calls and tools. Specifically, compound AI systems [19, 29] focus on the systemic perspective (the interconnection of components) and overlook the multi-step aspect (agent vs workflow). Meanwhile, the literature in language agents [25] neglects the efficiency of LM pipelines (also called workflows) [20] that leverage the task structure to reduce the burden in decision-making and supervise the LM with clearer instructions.
As shown in Fig. 2, we organize this survey as follows:
  1. a)
    Sec. 2 positions this survey in the broader literature of scaffolded LMs and recent breakthroughs in parametric training with RL.
  2. b)
@@
    Sec. 3 focuses on how the post-training interface of LMs interacts with the scaffold. We identify two types of scaffold—agents and workflows—and focus on important design aspects of each. Finally, we reflect on the trend of recent benchmarks and highlight the ability of scaffolded LMs for mixed-autonomy, a setting where human and AI share control and decision-making.
  3. c)
    Sec. 4 organizes the intricate literature that uses an LM as an optimizer for non-parametric training. We categorize these methods into whether they train LMs, agents, or workflows.
  4. d)
    Sec. 5 discusses a key capability missing from current scaffolded LMs: the ability to inhabit streams of experience [2, 30] and describes opportunities with non-parametric training.
Additionally, to help readers navigate this survey, we provide a glossary in Fig. 3.
Training of Scaffolded Language Models With Language Supervision
@@
Related Works §2 Autonomous Agents §2.1 Parametric Training §2.2 Scaffolded Language Models §3 Post-Trained Language Models §3.1 Tools §3.2 Multi-step Process §3.3 Agent §3.3.1 Workflow §3.3.2 Benchmarks §3.4 Training with Language Supervision §4 Prompt Optimization §4.1 Experiential Learning §4.2 Insights §4.2.1 Tools §4.2.2 AutoDiff frameworks §4.3 Graph Abstraction §4.3.1 Graph Execution and Optimization §4.3.2 Beyond Episodic Learning §5 Scalability §5.1 Plasticity and Stability §5.2 Efficiency §5.3 Interpretability §5.4
@@
Figure 2: Organization of this survey. This survey introduces a unifying paradigm centered on learning from language using the non-parametric variables of a scaffolded language model. Despite the complexity and fragmentation of the literature across domains such as language agents, prompt optimization, experiential learning, and compound AI systems, we adopt a deliberately simple and clear organizational structure. This simplicity is not a limitation, but a reflection of our effort to distill a coherent framework from a rapidly evolving field. It enables us to outline key research opportunities in learning with non-parametric variables in mixed-autonomy settings.
Term Definition
Scaffold (See Sec. 3) A surrounding structured code that manages control flow (multi-step process), maintains chat history, and executes tools based on responses of the LM.
Chat format (See Sec. 3.1) A chat format is a protocol that turns a structured conversation into a single sequence of strings that a post-trained LM both expects as input and emits as output.
Tool (See Sec. 3.2) An external capability the model can invoke via the Tools API (e.g., developer-defined functions, search engine). A function is a specific tool defined by a JSON schema. Non-parametric training is used to implement or modify functions.
Tool call (See Sec. 3.2) A specific type of response from the LM that includes a call to one of the available tools and a corresponding plain text for that tool (e.g., a query for a search engine); for a function call, it corresponds to a structured JSON.
Mixed-autonomy (See Sec. 3.4) A setting where humans and AI work together, sharing control and decision-making responsibilities to achieve a goal. Typically, the human delegates control to the AI when appropriate, dividing the task according to their respective strengths. For instance, in vibe coding, the human guides intent and judges quality, while the AI proposes implementations and refactors.
Prompt (See Sec. 4) The prompt includes the system and user message that is provided to an LM at the start of the chat. The user message specifies the task and data for the episode. The system message is shared across tasks, it sets the system behavior by specifying available tools, role of the scaffolded LM, high-level instructions, constraints, and any other information required to perform the task. Non-parametric training is used to improve the system prompt, e.g., by adding high-level insights or improving tool description.
Non-parametric variables (See Sec. 4) Prompts and tools of a scaffolded LM. Agents have a single prompt while workflows use a different prompt at each step. Non-parametric variables are expressed either in language or code.
LM optimizer (See Sec. 4) An LM optimizer prompts an LM with an objective function expressed in natural language to generate updated non-parametric variables.
Language supervision (See Sec. 4) It refers to the objective function and execution feedback, both expressed in natural language, that are provided to the LM optimizer.
Episode (See Sec. 5) Each episode begins with a new chat history when the human delegates control to the scaffolded LM via an instruction and concludes once the scaffolded LM successfully completes the task or the human resumes control. Importantly, during an episode, the human can offer corrections or provide guidance. Its textual form corresponds to an execution trace and feedback, which is used to prompt the LM optimizer.
Figure 3: Glossary of core terms.
2 Related Works
This section clarifies the position of this survey in the broader literature of autonomous agents [31, 25, 32] and parametric training [33].
2.1 Autonomous Agents
In this survey, we deliberately adopt the broader scaffold abstraction that encompasses both agents and pre-structured workflows. This is unlike existing surveys that focus on agents as the primary abstraction [31, 25, 32]. In particular, they focus on systematizing components such as perception, planning, action loops, tool use, memory, and environment interaction [31, 25], or feedback mechanisms (e.g., human preferences) to improve agent behavior [32]. We separate learning paradigms into (i) parametric training that updates LM weights (e.g., supervised learning, RL) and (ii) non-parametric training that updates non-parametric variables including the prompt and tools based on language-feedback. We argue that this separation is crucial for mixed-autonomy settings: non-parametric updates are interpretable and easily supervised by humans at high frequency, and they synergize with parametric post-training to deliver data-efficient learning. Using this taxonomy, we unify prompt optimization, experiential learning, and AutoDiff-style frameworks under a single lens.
2.2 Parametric Training
The discussion of parametric training [34, 35, 36] is important to understand the relevance of non-parametric training in the context of recent breakthrough in post-training of LMs with RL [33, 35, 34]. We start with an example in prompt optimization where recent methods that use parametric training to learn to reason offer much better performance [33]. To understand this, we compare RL with verifiable reward [33] with previous approaches. Provided with a reward that measures the performance of a model on a downstream task by verifying that the model’s answer match the ground-truth response [37]:
$r\triangleq\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}\Big[\mathbb{E}_{\mathbf{y}\sim p(\mathbf{x})}[\mathbbm{1}\{\mathbf{y}=\mathbf{y}^{\star}(\mathbf{x})\}]\Big],$ (1)
previously the following approaches have been used. One approach is to use an LM optimizer to generate a prompt and evaluate it against the reward. These methods use search methods such as MCTS [38] or beam search [39]. Another approach train the LM optimizer with RL to generate better prompts [40, 41, 42, 43]. Similarly, another line of work [44] proposes test-time editing of the instruction, resulting in a query-dependent LM optimizer. Tempera [44] uses a restricted action space (e.g., modify the verbalizer) to reduce variance in RL training. They define the reward as the relative performance before and after the edit on the downstream task. Recently, with stronger base models, LMs can directly optimize Eq. 1 with RL [33]. Importantly, doing so results in strong generalization, improvements beyond the domain at hand [45, 46, 47] and opens up a new axis for scaling, test-time compute. Similar examples can be found in scaffolded LMs trained with RL. SWE-RL [47] compares the generated patch with the GT patch on SWE-Bench. They do not use test cases as the reward signal, which can lead to reward hacking [48]. Specifically, SWE-RL is based on the Agentless [49] and applies RL on the patch generation step. They find that, unlike SFT, training with RL on SWE-Bench [50] leads to improvement across domains, including code generation with library use [51], code reasoning [52], math [53], and language understanding [54]. Thus, for tasks that are easily verifiable, parametric training works very well [55]. However, as we discuss in Sec. 5, non-parametric training offers interesting avenues of research in circumventing catastrophic forgetting and the plasticity-stability dilemma, allowing continuous learning across episodes [2].
@@ 3 Scaffolded Language Model
A scaffolded LM refers to an LM embedded in a multi-step process with tools. Such a structure allows for extending an LM’s utility to interact with digital systems in similar ways as humans do [5]. Scaffolded LMs involve tools and thus differs from meta-generation techniques [56] that employ multiple calls to an LM [57, 58, 59]. Meta-generation mimics search algorithms [60] for reasoning tasks [37] where each step consists of a call to an LM. Importantly, in scaffolded LMs, each call to an LM expects a tool response.
To provide readers with a practical understanding of how this structure allows LMs to perform sophisticated cognitive tasks and improve from non-parametric training, we organize this section as follows. We introduce the post-training interface of language models that allows interaction with tools and defines their behavior in natural language. Next, we discuss multiple aspects of tools including compositionality, reliability, safety, and built-in tools. Furthermore, we show how the multi-step process can be controlled to implement priors into how the task is solved. Finally, we discuss benchmarks to understand the shift induced by scaffolded LMs. Specifically, these benchmarks highlight the omnipresence of scaffolded LMs in mixed-autonomy settings, where the design of the scaffold defines the interaction dynamics between human and AI agents for sharing control and decision-making.